Daily Report

CPUC/ANF Environmental Monitoring Activities

Pacific Pipeline Construction Project

Aspen Environmental Group Monday, January 11, 1999

The following is a brief summary of the environmental monitoring activities that were completed between January 11, 1999 1:00 and 1:00 on January 12, 1999. This summary is based on reports directly received from the field and has not been edited. If you have any questions please contact (818) 597-3407. Please note the item number and the correction that should be made. This report is intended to serve as a communication document and as an instrument for quality control.

Construction Highlights

01/11/99 (Monday)

Segment 1 (Emidio Pump Station to Copco Road): Report for 11 January 1999: Backfill and clean up occurred again today at the Warbler Skip, ~MP 21.57. Work on Tejon Bench was limited to a grading crew and coating near the I-5/Tejon bore, to the east. Welders, x-ray, and coating worked on the tie in to the I-5/Tejon bore on the west side. Just below them, padding and FOC stringing were done, with minor backfill and clean up as well. The paving crew did some work along the I-5 frontage road, then came back to work on Grapevine Road. Continuing calibration of equipment was performed at Grapevine Station. The crew at Grapevine Creek, ~MP 15.35, backfilled to the north and prepared for testing on the south. Welders and x-ray worked on pipe above ground on the north side of the California Aqueduct crossing, MP 12.22. Surplus pipe and pups continued to be welded along Highway 166, just east of Emidio Pump Station. More 8" pipe was also welded for the future dewatering into Stream Crossing #1. The hydrotest crew began filling more water into the pipe from Emidio.

Segment 3 (Sylmar to Los Angeles) Report for 11 January 1999: Electrical crews worked at Cornfield Yard, Balboa Blvd, and Highland Avenue. Clean-up and recontouring occurred sporadically along the ROW.

Segment 5 (Lynwood to Long Beach): Report for 1-11-99. The bore crew continued work digging under Alameda St at the Edison station site. In addition, crews inserted several 29'' diameter sleeves and anticipate completion of the bore by Tuesday. Meanwhile, crews working nearby progressed with the clean up and grading of the rail road right of way. Further, crews removed debris and equipment from the right of way at Auto Center Drive. At the Tosco refinery crews continued wrapping and installing oil overflow drain lines to the sump pump while work was conducted on the station electrical system. Similarly, the station commission crew was on site to begin final mechanical and electrical inspection of the metering system. While work was being conducted within Tosco a jet fuel leak was discovered along the eastern fence adjacent to the PPSI bore pit which connects Tosco to the main section of the PPSI line. No work in this section was currently in progress although a general clean up of the right of way was conducted on 1-8-99. Emergency crews were notified and a berm was placed around the approximately 80 x 20 yard spill. Local and refinery response teams in addition to police, fire and Haz Mat units were on site. Foam was applied to the spill and pump trucks were utilized to remove the fuel. Further, crews anticipate removing contaminated soil throughout the evening. Segment 4B (Lynwood to El Segundo): Crews continued brick work at the Lynwood station. In addition, gravel was spread within the station while crews progressed hanging metal doors and removing excess spoils from the right of way. Crews working at the El Segundo station continued final electrical work and began removing spoils from the station site.

Inspection Summar

In the past 24 hours, 3 inspection event(s) were reported for 3 spread(s). Of these inspections, 0 were non-compliance event(s). Also 0 non-compliance resolution report(s) was resolved.

Non-Compliance Reports (NCR's) Issued

No new non-compliance reports were issued during the last reporting period.

Non-Compliance Resolution Reports (NCRR's) Issued

No new non-compliance reports were issued during the last reporting period.

Summary of Follow-up Items for Today

No follow-up items were reported for today.

Back to Daily Reports

Back to MMCRP Home Page